# Dyn-HaMR: Recovering 4D Interacting Hand Motion from a Dynamic Camera

### v2

@@ 1 Introduction
  • Proposing a data-driven hand motion prior combined with biomechanical constraints, allowing realistic and complex hand interactions to guide the optimization.
@@ 2 Related Work
Recently, pose and shape estimation of two hands, which we refer as bimanual has rapidly progressed. Among them, [35] has further scaled up the single pose estimation. Li et al. [23] developed a mesh regression network based on GCN that employs pyramid features and learned implicit attention between two hands. Moon el at. [30] proposed to improve in-the-wild hand reconstruction accuracy with a strategy to bridge the domain gap between multi-camera datasets and in-the-wild datasets. Ren et al. [39] used a variational autoencoder (VAE) as a prior for interacting hand reconstruction. Yu et al. [51] introduced a one-stage hand reconstruction pipeline for two hands by an attention aggregation mechanism with a 2D Gaussian heatmap and cross-hand attention. All these methods suffer from severe depth ambiguity and increased jitter in reconstruction due to the coupling of camera and hand poses. Moreover, their reliance on weak-perspective camera models inherently limits their ability to perform effectively in dynamic camera scenarios. In contrast, our optimization-based approach can reconstruct 4D global motion of two hands from complex scenes captured by a moving camera, while more accurately modelling their interaction using a learned hand motion prior.
@@ 3.1 Hand Tracking and Hierarchical Initialization
The first term aligns the reprojection of ${}^{\mathrm{w}}\mathbf{Q}^{h}$ with the initial 2D observations $\{\hat{\mathbf{j}^{h}_{t}}\}$:
@@
Here, $\tilde{\bm{J}_{t}^{h}}=\Pi(^{\mathrm{w}}\mathbf{J}_{t}^{h},\mathbf{R}_{t},\omega,\bm{{\tau^{\mathrm{c}}}}_{t},\mathbf{K})$ and $\Pi$ is the perspective camera projection with camera intrinsics $\mathbf{K}\in\mathbb{R}^{3\times 3}$. $\mathbf{C}^{h}_{t}$ is a mask obtained from the joint visibility and $\tilde{\mathbf{j}_{t}^{h}}$ is the reprojected 2D keypoints from the current 3D keypoints ${}^{\mathrm{w}}\mathbf{j}_{t}^{h}$. $\rho(\cdot)$ is the Geman-McClure robust function [11].
@@ 4 Experimental Evaluation
Among these, the most suitable datasets for our task are H2O and HOI4D, as they contain videos recorded by dynamic cameras with large and frequent displacement and the camera poses are provided. For all of the datasets used, we follow the official split to conduct our experiments.
@@
We compare our method against state-of-the-art approaches of ACR [51], IntagHand [23] and HaMeR [35] in terms of monocular hand reconstruction and global hand motion estimation. We provide further details on these methods in our suppl. material.
@@ 4.1 Results
Figure 4: Qualitative comparison with state-of-the-art method HaMeR [35]. It can be seen that our method recovers significantly more plausible global hand motion. first row is from H2O dataset [19], while second & third rows are from HOI4D dataset [27].
@@
Tabs. 2 and 3 (as well as our suppl. material) reports our quantitative results on recovering local and global 4D motion under the presented metrics, using the official test split for the egocentric dynamic camera views. It can be observed that our Dyn-HaMR outperforms all of the state-of-the-art methods [51, 35, 23] by a large margin in terms of G-MPJPE and GA-MPJPE. This is due to the superiority of our global 4D interacting hand motion recovery.
@@
Fig. 4 presents a qualitative comparison against state-of-the-art methods, revealing significant improvements in generating realistic and plausible motion trajectories, particularly with enhanced depth reasoning under inter-hand occlusions. It can be seen that our method can better recover the actual global trajectories and respect the motion of both the hands and the moving camera, whilst other methods suffer from depth ambiguity, being agnostic to camera pose & displacements. The challenging entanglement of hand and camera motion would further degenerate the scale deduction and lead to implausible motions. We present further qualitative results in motion reconstruction on HOI4D, FPHA and EgoDexter datasets in the suppl. material. Our suppl. video provides in-the-wild acquisition of hand interactions from dynamic cameras.
Table 2: Quantitative evaluation results for H2O dataset. Our method demonstrates significant improvements over state-of-the-art approaches in recovering both local and global 4D hand motion, with additional gains achieved when incorporating Stage III.
@@
Ablation studies
To fully assess the effectiveness of Dyn-HaMR, we perform further ablation studies on the pipeline design analyzing the contribution of each component. In particular, we investigate the effectiveness of the key components: (i) $\mathcal{L}_{bio}$, (ii) $\mathcal{L}_{pen}$, (iii) the interacting hand motion prior module in Stage III, and (iv) generative infilling, where we replace the generative infilling module with a simple interpolation. It can be seen from Tabs. 2 and 4 that incorporating Stage III can boost the performance by a considerable margin as it provides a well-learned motion prior information for the final stage optimization and yields more plausible and smoother 4D trajectory reconstructions. Moreover, as shown in Fig. 3, adding the $\mathcal{L}_{bio}$ and $\mathcal{L}_{pen}$ can significantly improve the motion quality making it more plausible and realistic.
Table 4: Ablation of pipeline components on H2O dataset. It shows the impact of removing different components from the pipeline on various performance metrics.
@@ 5 Conclusion
While Dyn-HaMR successfully works for limited time-horizons, extending it to long sequences with generative, extrapolation capabilities remains to be explored. Developing a regression-based method is one of the possible directions to deal with long sequences. We will also work on improving the hand priors and incorporate object interactions.
@@
This document supplements our main paper in terms of implementation details and additional qualitative and quantitative evaluations. Moreover, we refer the reader to our supplementary video for the best view of the resulting motions and for a more comprehensive exposition, where we show promising results on in-the-wild hand reconstruction and qualitative comparisons against the state-of-the-art interacting hand reconstruction methods [35] under the challenging dynamic camera scenarios.
Appendix A Implementation Details and Data Processing
@@ A.1 Initializing Motion States in Camera Frame
First, we employ a 2D hand pose estimation model, ViTPose, following [48, 35], known for its performance in hand detection and palm localization. Despite its strength in detecting global hand regions, the model often produces jittery and inaccurate joint positions, making it insufficient for subsequent optimization processes. As a remedy, we refine the 2D inputs by cropping the image based on bounding boxes calculated from ViTPose’s 2D keypoint predictions. Specifically, for a set of keypoints ${}^{\mathrm{vit}}\hat{\bm{J}}^{h}_{t}$, the bounding box is calculated by the point sets with a confidence filter $\epsilon_{b}=0.5$ and an extension coefficient of $200\%$. To initialize the MANO parameters $\{\bm{\theta}^{h}_{t},\bm{\beta}^{h}_{t},^{\mathrm{c}}\bm{\phi}^{h}_{t}\}$, we run the state-of-the-art hand reconstruction method of [35], using the officially released checkpoints, on the image patches restricted to the calculated bounding box. To estimate translation $\bm{{}^{\mathrm{c}}\mathrm{\tau}}^{h}_{t}=(x,y,d)$ of the two hands in 3D space, where $d$ is the direction along depth, we simulate various versions of $\bm{{}^{\mathrm{c}}\mathrm{\tau}}^{h}_{t}$ in the camera coordinate system based on the predicted weak-perspective camera parameters $(s,t_{x},t{y})$, assuming a fixed focal length $f=1000$ following [35]:
$x=t_{x},y=t_{y},d=\frac{2f}{s\times s_{\mathbf{I}}}.$ (14)
Here $s_{\mathbf{I}}$ is the image size. Alternatively, the camera translation in the camera coordinate system can also be acquired by solving the PnP algorithm (i.e. RANSAC [9]) with the 3D keypoints ${}^{\mathrm{c}}\bm{J}^{h}_{t}$ and their corresponding 2D projections $\hat{\bm{J}}^{h}_{t}$ on the image plane. Finally, we infill the missed frames if the interval is less than 50 frames using the approach described in Sec. 3.1.
@@ A.2 Optimization Scheme
Appendix B Additional Experiments
@@
B.1 Results
Figure 6: Comparison of global trajectory on HOI4D [27].
@@ B.1 Results
Local motion estimation
Global motion estimation
As introduced in Sec. 4.1 of the main paper, H2O [19] and HOI4D [27] contain dynamic camera videos with available camera poses to convert the hand poses from the camera coordinate system to the world coordinate system. To evaluate global motion recovery performance, we have conducted qualitative evaluations on the aforementioned four egocentric hand-object interaction datasets mentioned in Sec. 4.1 as well as on in-the-wild videos. In this section, we first present qualitative results on these datasets shown in Figs. 9, 11 and 10, where our method produces plausible 4D global motion with trajectories in the world coordinate system, while previous state-of-the-art methods fail to capture the global motion in 3D space, especially from dynamic cameras. Furthermore, our method yields more plausible depth reasoning in the bimanual setting and significantly reduces the jitter in translations. To quantify the reconstruction accuracy and errors, we evaluate the Translation Error and Jerk in Tab. 6, where we can observe significant improvements over the state-of-the-art approaches. Specifically, we evaluate the RTE score for each of the motion trajectories in the world coordinate system. It can be observed that our method consistently archives the lowest translation error across datasets. We also conduct further analysis on the HOI4D dataset as shown in Fig. 6, which contains large camera displacements. Notably, we achieve 3.89% in Trans Err against 18.98% of HaMeR [35] on this dataset. We encourage to browse the reconstruction results in our supplementary videos, which clearly demonstrates the superiority of our pipeline against the state-of-the-art methods.

### v3

@@ 1 Introduction
  • Leveraging a data-driven hand motion prior with biomechanical constraints to guide the optimization, enabling the recovery of realistic and complex hand interactions
@@ 2 Related Work
Pose and shape estimation of two hands (bimanual) has rapidly progressed recently [38, 26, 33, 42, 54, 24]. Among them, [38] scaled up the single pose estimation. Li et al. [26] developed a mesh regression network based on GCN that employs pyramid features and learned implicit attention between two hands. Moon el at. [33] proposed to improve in-the-wild hand reconstruction accuracy with a strategy to bridge the domain gap between multi-camera datasets and in-the-wild datasets. Ren et al. [42] used a variational autoencoder (VAE) as a prior for interacting hand reconstruction. Yu et al. [54] introduced a one-stage hand reconstruction pipeline for two hands by an attention aggregation mechanism with a 2D Gaussian heatmap and cross-hand attention. SignAvatars [55] used a multi-stage optimization pipeline to reconstruct 3D sign actors. All these methods suffer from depth ambiguity and increased jitter in reconstruction due to the coupling of camera and hand poses. Moreover, reliance on weak-perspective camera models inherently limits their ability to perform effectively in dynamic scenarios. In contrast, our optimization-based approach can reconstruct 4D global motion of two hands from complex scenes captured by a moving camera, while more accurately modelling their interaction using a learned hand motion prior.
@@ 3.1 Hand Tracking and Hierarchical Initialization
The first term aligns the reprojection of ${}^{\mathrm{w}}\mathbf{Q}^{h}$ with the initial 2D observations $\hat{\mathbf{J}^{h}_{t}}$:
@@
Here, $\tilde{\mathbf{J}_{t}^{h}}=\Pi(^{\mathrm{w}}\mathbf{J}_{t}^{h},\mathbf{R}_{t},\omega,\bm{{\tau^{\mathrm{c}}}}_{t},\mathbf{K})$ and $\Pi$ is the perspective camera projection with camera intrinsics $\mathbf{K}\in\mathbb{R}^{3\times 3}$. $\mathbf{C}^{h}_{t}$ is a mask obtained from the joint visibility and $\tilde{\mathbf{J}_{t}^{h}}$ is the reprojected 2D keypoints from the current 3D keypoints ${}^{\mathrm{w}}\mathbf{J}_{t}^{h}$. $\rho(\cdot)$ is the Geman-McClure robust function [12].
@@ 4 Experimental Evaluation
  • Ego-Exo4D [14] is large-scale multi-view dataset with multi-modal annotations capturing simultaneous ego and multiple exo videos. We evaluated on the subset of hand ego-pose benchmark.
  • HOT3D [3] is an egocentric dataset with multiview synchronized image streams recorded with two head-mounted devices, which provides accurate 3D poses and shapes of hands and objects.
Among them, we use the validation set for HOT3D and Ego-Ego4D datasets since their test set annotations (required to obtain GT global motion) are not publicly available this point. For other datasets used, we follow the official split to conduct our experiments.
@@
We compare our method against state-of-the-art approaches of ACR [54], IntagHand [26] and HaMeR [38] in terms of monocular hand reconstruction and global hand motion estimation. We initialize Ours (Dyn-HaMR) from [38] to get the results in the tables.
@@ 4.1 Results
Figure 4: Qualitative comparison with state-of-the-art method HaMeR [38]. It can be seen that our method recovers significantly more plausible global hand motion. first row is from H2O dataset [22], while the second row is from HOI4D dataset [30].
@@
Tabs. 2 and 3 (as well as our suppl. material) reports our quantitative results on recovering local and global 4D motion under the presented metrics, using the official test split for the egocentric dynamic camera views. It can be observed that our Dyn-HaMR outperforms all of the state-of-the-art methods [54, 38, 26] by a large margin in terms of G-MPJPE and GA-MPJPE.
@@
Fig. 4 presents a qualitative comparison against state-of-the-art methods, revealing significant improvements in generating realistic and plausible motion trajectories, particularly with enhanced depth reasoning under inter-hand occlusions. Our method can better recover the actual global trajectories and respect the motion of both the hands and the moving camera, whilst other methods suffer from depth ambiguity, being agnostic to camera pose & displacements. The entanglement of hand and camera motion further degenerates the scale estimation and lead to implausible motions. Moreover, Tab. 4 and Tab. 5 show the results on challenging two-hand object interaction scenarios, where further compare our method with the canonical HaMeR + DPVO demonstrating the vitality of our optimization. We present further qualitative results in motion reconstruction on HOI4D, FPHA and EgoDexter datasets in the suppl. material. Our suppl. video provides in-the-wild capture of hand interactions from dynamic cameras.
Table 2: Quantitative evaluation results for H2O [22] dataset. Our method demonstrates significant improvements over state-of-the-art approaches in recovering both local and global 4D hand motion, with additional gains achieved incorporating Stage III.
@@
Ablation studies
To fully assess the effectiveness of Dyn-HaMR, we perform further ablation studies on the pipeline design analyzing the contribution of each component. In particular, we investigate the effectiveness of the key components: (i) $\mathcal{L}_{bio}$, (ii) $\mathcal{L}_{pen}$, (iii) the interacting hand motion prior module in Stage III, and (iv) generative infilling, where we replace the generative infilling module with a simple interpolation. It can be seen from Tabs. 2 and 6 that incorporating Stage III can boost the performance by a considerable margin as it provides a well-learned motion prior information for the final stage optimization and yields more plausible and smoother 4D trajectory reconstructions. Moreover, as shown in Fig. 3, adding the $\mathcal{L}_{bio}$ and $\mathcal{L}_{pen}$ can significantly improve the motion quality making it more plausible and realistic.
@@
Table 4: Quantitative evaluation Results on Ego-Exo4D [14] dataset. We only report the positional error metrics as the GT annotation is too discrete to extract a meaningful acc.
Method Jerk $\downarrow$ PA-MPJPE $\downarrow$ G-MPJPE $\downarrow$
ACR 245.12 19.45 291.34
IntagHand 175.65 24.32 275.89
HaMeR 195.45 15.87 267.75
HaMeR + DPVO 200.45 15.87 213.75
Ours (Dyn-HaMR) 5.26 14.34 53.89
Table 5: Qualitative results on HOT3D [3] dataset. We compare our method with the canonical HaMeR + DPVO baseline.
Method Jerk $\downarrow$ PA-MPJPE $\downarrow$ G-MPJPE $\downarrow$ Acc Err $\downarrow$
ACR 153.45 16.41 159.34 16.45
IntagHand 171.24 21.75 165.42 15.12
HaMeR 189.62 10.43 155.98 13.78
HaMeR + DPVO 195.77 10.43 129.45 12.78
Ours (Dyn-HaMR) 4.18 8.87 42.36 4.95
Table 6: Ablation of pipeline components on H2O [22] dataset. It shows the impact of removing different components from the pipeline on various performance metrics.
@@ 5 Conclusion
While Dyn-HaMR successfully works for limited time-horizons, extending it to long sequences with generative, extrapolation capabilities remains to be explored. Developing a regression-based method is one of the possible directions to deal with long sequences. We will also work on improving the hand priors extending [17] and incorporate object interactions.
Acknowledgments
T. Birdal and S. Zafeiriou acknowledge support from the Engineering and Physical Sciences Research Council [grant EP/X011364/1]. T. Birdal was supported by a UKRI Future Leaders Fellowship [grant number MR/Y018818/1] as well as a Royal Society Research Grant RG/R1/241402.
@@ Acknowledgments
The following sections supplement our main paper in terms of implementation details and additional qualitative and quantitative evaluations. Moreover, we refer the reader to our supplementary videos on the project webpage to better perceive the resulting motions and for a more comprehensive exposition, where we show promising results on in-the-wild hand reconstruction and qualitative comparisons against the state-of-the-art interacting hand reconstruction methods [38] under the challenging dynamic camera scenarios.
Appendix A Implementation Details and Data Pre-Processing
@@ A.1 Initializing Motion States in Camera Frame
First, we employ a 2D hand pose estimation model, ViTPose, following [51, 38], known for its performance in hand detection and palm localization. Despite its strength in detecting global hand regions, the model often produces jittery and inaccurate joint positions, making it insufficient for subsequent optimization processes. As a remedy, we refine the 2D inputs by cropping the image based on bounding boxes calculated from ViTPose’s 2D keypoint predictions. Specifically, for a set of keypoints ${}^{\mathrm{vit}}\hat{\bm{J}}^{h}_{t}$, the bounding box is calculated by the point sets with a confidence filter $\epsilon_{b}=0.5$ and an extension coefficient of $200\%$. To initialize the MANO parameters $\{\bm{\theta}^{h}_{t},\bm{\beta}^{h}_{t},^{\mathrm{c}}\bm{\phi}^{h}_{t}\}$, we run the state-of-the-art hand reconstruction method of [38], using the officially released checkpoints, on the image patches restricted to the calculated bounding box. To estimate translation $\bm{{}^{\mathrm{c}}\mathrm{\tau}}^{h}_{t}=(x,y,d)$ of the two hands in 3D space, where $d$ is the direction along depth, we simulate various versions of $\bm{{}^{\mathrm{c}}\mathrm{\tau}}^{h}_{t}$ in the camera coordinate system based on the predicted weak-perspective camera parameters $(s,t_{x},t{y})$, assuming a fixed focal length $f=1000$ following [38] using $x=t_{x},y=t_{y},d=\frac{2f}{s\times s_{\mathbf{I}}}$, where $s_{\mathbf{I}}$ is the image size. Alternatively, the camera translation in the camera coordinate system can also be acquired by solving the PnP algorithm (i.e. RANSAC [10]) with the 3D keypoints ${}^{\mathrm{c}}\bm{J}^{h}_{t}$ and their corresponding 2D projections $\hat{\bm{J}}^{h}_{t}$ on the image plane. Finally, we infill the missed frames if the interval is less than 50 frames using the approach described in Sec. 3.1.
@@
Handling hallucinations
HaMeR can yield erroneous hallucinations – such as multiple hands in the same location, incorrect handedness, or implausible poses. Specifically, HaMeR’s detector can produce overlapping bounding boxes (bboxes) without suppression, leading to redundant or inconsistent predictions as both ViTPose and HaMeR process each bbox independently. As a remedy, we use ViTPose to extract 2D keypoints with confidences (IoU $>0.9$) and instead of processing all overlapping detections, we retain only the bbox with the highest confidence before feeding it into HaMeR. We also filter any detection that appears in $<$10 frames, reducing false positives. Incorrect handedness, where one hand is occasionally confused as its opposite, can be identified by the sudden change in the bbox (IoU with the previous bbox $<$0.1), allowing us to mark these frames as invalid prior to generative infilling.
A.2 Optimization Scheme
@@
Appendix B Evaluation Metrics
To compute G-MPJPE and GA-MPJPE, we first align the first two frames (G-MPJPE) or the whole sequence (GA-MPJPE) of hand motion with the GT using Umeyama method. We then transform the prediction to align with the GT before computing the MPJPE as the mean $L_{2}$ distance between each predicted and GT joint. To compute the Acc Err, we followed the common practice in HMP, GLAMR and SLAHMR ignoring the division, $\alpha_{i}^{t}=v^{t-1}_{i}-2v^{t}_{i}+v^{t+1}_{i}$, where $\alpha_{i}^{t}$, $v^{t-1}_{i}$ are the acc. and velocity at timestep $t$ without the division of discretized time step. We further analyze the effect of the different computation of acceleration where the discretized time step is applied or not. We report both Acc Err with and w/o division ($mm^{2}$ and $m/s^{2}$) in Tab. 7, where we can observe that our method keeps consistently better results under different computation method.
Table 7: Acceleration analysis on HOT3D dataset. Acc Err is reported w/o the div. of $\omega^{2}$ (left) in $mm/s^{2}$ and with the div. of $\omega^{2}$ in $m/s^{2}$ (right). Lower ($\downarrow$) is always better.
Method Acc Err (w/o) $\downarrow$ Acc Err (with) $\downarrow$
ACR [54] 16.45 14.82
IntagHand [26] 15.12 13.62
HaMeR [61] 13.78 12.41
HaMeR + DPVO [50] 12.78 11.51
Ours (Dyn-HaMR) 4.95 4.46
[…]
Figure 6: Comparison of global trajectory on HOI4D [30].
Appendix C Additional Experiments
@@
Figure 7: Qualitative comparison with state-of-the-art method HaMeR [61] on in-the-wild online videos.
C.1 Results
Local motion estimation
Global motion estimation
As introduced in Sec. 4.1 of the main paper, H2O [22] and HOI4D [30] contain dynamic camera videos with available camera poses to convert the hand poses from the camera coordinate system to the world coordinate system. To evaluate global motion recovery performance, we have conducted qualitative evaluations on the aforementioned four egocentric hand-object interaction datasets mentioned in Sec. 4.1 as well as on in-the-wild videos. In this section, we first present qualitative results on these datasets shown in Figs. 10, 13 and 12, where our method produces plausible 4D global motion with trajectories in the world coordinate system, while previous state-of-the-art methods fail to capture the global motion in 3D space, especially from dynamic cameras. Furthermore, our method yields more plausible depth reasoning in the bimanual setting and significantly reduces the jitter in translations. To quantify the reconstruction accuracy and errors, we evaluate the Translation Error and Jerk in Tab. 9, where we can observe significant improvements over the state-of-the-art approaches. Specifically, we evaluate the RTE score for each of the motion trajectories in the world coordinate system. It can be observed that our method consistently archives the lowest translation error across datasets. We also conduct further analysis on the HOI4D dataset as shown in Fig. 6, which contains large camera displacements. Notably, we achieve 3.89% in Trans Err against 18.98% of HaMeR [38] on this dataset. We encourage to browse the reconstruction results in our supplementary videos, which clearly demonstrates the superiority of our pipeline against the state-of-the-art methods.
@@ C.1 Results
Figure 11: In-the-wild 4D global hand motion reconstructions on HOT3D [3] (top and bottom rows) and Ego-Exo4D [14] (middle row) datasets.